Can you tell me how to do this problem. I have not had Algebra before so I need to be sure I'm on the right track.
2x + 6 =30

Answer by Tias_411(14)

first you want to get your (2x) by itself so you subtract 6 from both sides of your equation. 2x + 6 =30
- 6 -6
___________
2x =24
*your 6's will cancel out.
now you want x by itself so you divide both sides by 2.
2x =24
/2 /2
x= 12
*your 2's canceled to get you x by itself.